# Absolute continuity of non-Gaussian and Gaussian multiplicative chaos measures

### v1

@@ Abstract.
In this article, we consider the multiplicative chaos measure associated to the log-correlated random Fourier series, or random wave model, with i.i.d. coefficients taken from a general class of distributions. This measure was shown to be non-degenerate when the inverse temperature is subcritical by Junnila (Int. Math. Res. Not. 2020 (2020), no. 19, 6169-6196). When the coefficients are Gaussian, this measure is an example of a Gaussian multiplicative chaos (GMC), a well-studied universal object in the study of log-correlated fields. In the case of non-Gaussian coefficients, the resulting chaos is not a GMC in general. However, for inverse temperature inside the $L^{1}$-regime, we construct a coupling between the non-Gaussian multiplicative chaos measure and a GMC such that the two are almost surely mutually absolutely continuous.
@@ 1. Introduction
This article considers the multiplicative chaos measures associated to non-Gaussian log-correlated fields and considers their shared properties with respect to Gaussian multiplicative chaos (GMC), formalized through the relation of absolute continuity. The model we consider is the log-correlated random Fourier series, or random wave model. We focus on the space $[0,1]$ for ease of exposition and address general bounded domains in $\mathbb{R}^{d}$, $d\geq 1$, in Section 4.3. We begin by defining the random Fourier series and stating the main theorem. After this, we review GMC, discuss related literature in the non-Gaussian setting, and provide an outline of the proof of the main theorem.
@@
$\mu_{n,\gamma,a}(\mathrm{d}t):=\frac{e^{\gamma S_{a,n}(t)}}{Z_{n,\gamma,a}(t)}\,\mathrm{d}t\,,$ (1.3)
where $Z_{n,\gamma,a}(t):=\mathbb{E}\left[e^{\gamma S_{n,a}(t)}\right]$ so that $\mu_{n,\gamma}[A]$ is a martingale for all Borel sets $A$.
In [Junnila], Junnila showed that, for $\gamma\in(0,\sqrt{2})$, there exists a non-degenerate measure $\mu_{\gamma,a}$ on $[0,1]$ such that $\mu_{n,\gamma,a}$ converges in probability to $\mu_{\gamma,a}$, as $n\to\infty$, in the space of Radon measures on $[0,1]$ equipped with the topology of weak convergence (for $\gamma\geq\sqrt{2}$, the limiting measure is degenerate). We call $\mu_{\gamma,a}$ the multiplicative chaos measure associated to the fields $(S_{n,a})_{n\in\mathbb{N}}$.
When $a:=g$ is a sequence of i.i.d. standard Gaussians, the non-degeneracy of $\mu_{\gamma,g}$ has been known since Kahane [kahanegmc], and $\mu_{\gamma,g}$ is called Gaussian multiplicative chaos (GMC). GMC is a fractal random measure with fascinating properties and many important applications in modern probability; we review this object in Section 1.1 below. Here, we mention that GMC is a universal object expected to appear when exponentiating a field that exhibits asymptotically Gaussian behavior and logarithmically-decaying correlations.
While the field $(S_{n,a}(t))_{t\in[0,1]}$ is log-correlated (shown in [Junnila, Section 3.1]), the chaos measure $\mu_{\gamma,a}$ is not a GMC in general, due to the contribution of, for instance, the first (non-Gaussian) term in the sum defining $S_{n,a}(t)$. This begs the question: for general sequences $a$, what properties does $\mu_{\gamma,a}$ share with GMC?
@@
Let $\gamma\in(1,\sqrt{2})$ and take $(a^{(i)}_{k})_{i\in\{1,2\},k\in\mathbb{N}}$ i.i.d. satisfying (1.2). There exists a sequence of i.i.d. standard Gaussians $g:=(g_{k}^{(i)})_{i\in\{1,2\},k\in\mathbb{N}}$ coupled to a:=($a:=(a^{(i)}_{k})_{i\in\{1,2\},k\in\mathbb{N}}$ such that
$\mu_{\gamma,g}\ll\mu_{\gamma,a}\ll\mu_{\gamma,g}\qquad\text{almost surely.}$.
The $(g^{(i)}_{k})_{k\in\mathbb{N},i\in\{1,2\}}$ are defined in (2.23), which makes use of (2.3), (2.19) and Lemma 2.6. We expect 1.1 can assist in proving a variety of almost-sure properties of both $\mu_{\gamma,a}$ and the associated approximating sequence of non-Gaussian fields $(S_{n,a})_{n\in\mathbb{N}}$ shared by the GMC $\mu_{\gamma,g}$ and the Gaussian log-correlated fields $(S_{n,g})_{n\in\mathbb{N}}$, for $\gamma\in(1,\sqrt{2})$.
Remark 1.1 (Absence of the $L^{2}$-regime).
Our theorem does not address $\gamma\in(0,1]$; the region $\gamma\in(0,1)$ is the so-called $L^{2}$-regime of multiplicative chaos. This is the regime of $\gamma$ in which the measure has finite second moments and as such the situation is typically more tractable. The absence of the $L^{2}$-regime from our result is not a coincidence coming from the choice of the dimension $d=1$. In Theorem 4.1, we extend our result to random Fourier series in bounded domains in $\mathbb{R}^{d}$, where the result again only stands outside the $L^{2}$-regime, i.e., $\gamma\in(\sqrt{d},\sqrt{2d})$. It is of significant interest to the authors to either extend 1.1 to all subcritical $\gamma$, or to show that, in some example, non-Gaussian and Gaussian multiplicative chaos actually behave differently in the $L^{2}$-regime. We discuss why our proof requires $\gamma>1$ in Section 1.3.
Remark 1.2 (Improvement with matching third moment).
Under the additional assumption that $\mathbb{E}[(a_{1}^{(1)})^{3}]=0$ (e.g., take $a_{1}^{(1)}$ to have symmetric distribution), we can extend the result to $\gamma\in(\sqrt{2/3},\sqrt{2})$. It would be interesting to examine whether assuming more moments of $a_{1}^{(1)}$ match those of a Gaussian further extends the range of $\gamma$ for which mutual absolute continuity holds. We elaborate on this in Section 4.1.
@@ 1.1. Background on GMC
Gaussian multiplicative chaos was introduced by Kahane in [kahanegmc] as a mathematical model for energy dissipation in turbulence, making rigorous a program initiated by Mandelbrot in [mandelbrot-turbulence]. Therein, considering a locally compact metric space $(T,\rho)$ with reference (Radon) measure $\sigma$, Kahane gave a rigorous interpretation of the measure
@@
where $(X(x))_{x\in T}$ formally denotes a log-correlated Gaussian field with $\sigma$-positive covariance kernel
@@
for $g$ bounded and continuous. The kernel $R$ is of $\sigma$-positive type if there exists a sequence $(R_{i})_{i}$ of continuous, nonnegative, and positive definite kernels $R_{i}:T^{2}\to\mathbb{R}_{+}$ with
@@
Formally, (1.6) means that the Gaussian field $X$ can be approximated by the (pointwise-defined) Gaussian fields $X_{n}=\sum_{i=1}^{n}X_{i}$, where $X_{i}$ is a centered Gaussian with covariance $R_{i}$. The measures
$M_{n,\gamma}(\mathrm{d}x):=e^{\gamma X_{n}(x)-\frac{\gamma^{2}}{2}\operatorname{Var}[X_{n}(x)]}\ \sigma(\mathrm{d}x)$
form a martingale sequence, so $M_{\gamma}$ in (1.4) may be defined as its limit. An important property proven in [kahanegmc] is that $M_{\gamma}$ is independent of the sequence $(X_{n})_{n\in\mathbb{N}}$ approximating $X$, so that we can assosciate a unique measure $M_{\gamma}$, called GMC, to the log-correlated Gaussian field $(X(x))_{x\in T}$.
Removing the restriction (1.6) was achieved in several steps. First, [robert2010gaussian] defined $M_{\gamma}$ in the case that $R$ is a stationary positive-definite kernel of the form (1.5) using a mollification approach, which in particular does not rely on a martingale approximation. Later, [shamov2016gaussian] defined the GMC associated to general Gaussian fields using an approach based on random shifts. Finally, [BerestyckiSimplePath] simplified the mollification approach and gave a simple proof that the limiting measure is independent of the chosen mollification procedure. More on these approaches, as well as a rich theory on the behavior of GMC and several applications, can be found in [RV14, berestycki-powell].
There are various examples of non-Gaussian log-correlated fields producing a limiting GMC coming from random matrix theory; see, for instance, [webb15, berestycki2018random, CN19, NSW20, claeys2021much, bf22, kivamae24, junnila2024multiplicative, lambert-najnudel]. In [claeys2021much], conditions on the exponential moments of the field under which one has a GMC limit are developed. This criterion and the fact that GMC is supported on its thick points have been used in [lambert2020maximum, lambert2024law] to determine the first order of the maximum of non-Gaussian log-correlated fields, see also [peilen24].
@@ 1.2. Results on non-Gaussian multiplicative chaos
$\mu(\mathrm{d}t)\stackrel{{\scriptstyle d}}{{=}}Z_{0}\mathbf{1}_{[0,1/2]}(t)\mu_{0}(2\mathrm{d}t)+Z_{1}\mathbf{1}_{[1/2,1]}(t)\mu_{1}(2\mathrm{d}t-1)\,,$ (1.7)
@@
where $\mu_{0}$, $\mu_{1}$ are independent copies of $\mu$ and $Z_{0}$, $Z_{1}$ are non-Gaussian, i.i.d. with mean 1. Motivated by Eq. 1.7, [bacry2003log] studies multiplicative chaos measures for which the underlying field is given by the mass of certain cones under an infinitely divisible, independently scattered random measure. They show that, for their construction, a limiting measure exists and give a criterion for non-degeneracy and existence of moments. In [10.1214/12-AOP810], also motivated by Eq. 1.7, so-called $\star$-scale invariant random measures are studied, which, under mild assumptions, are formally of the form $e^{\gamma L_{x}}\;\mathrm{d}x$ with $L$ a log-correlated and infinitely divisible process. They show the existence of $\star$-scale invariant random measures, characterize them (up to technical assumptions), and give a condition for non-degeneracy.
@@
Brownian multiplicative chaos is the chaos measure one obtains by formally exponentiating the square-root of the occupation field of 2D Brownian motion. This object was constructed in [BBK94] in the $L^{2}$-regime and in [AHS20, Jeg20a] in the full sub-critical regime. Great progress has been made in its study and characterization since then [loop-soup, jego21, jego23].
@@
In [fan2023trigonometric], random trigonometric series of the form $\sum_{n}\alpha_{n}X_{n}\cos(nt+\Phi_{n})$ are considered, where $\alpha_{n}\in\ell^{4}(\mathbb{R})\setminus\ell^{2}(\mathbb{R})$, $\Phi_{n}$ are i.i.d. with $\Phi_{1}\sim\mathrm{Unif}([0,2\pi))$, and $X_{n}$ are independent with $\mathbb{E}[X_{n}^{2}]=1$ and $\mathbb{E}[X_{n}^{2m}]=O(K^{m}m!)$ for some $K>0$ and all $m\geq 2$, $n\geq 1$. They study the associated chaos measures and give results on the Hausdorff dimension. It is worth mentioning that in the case $\alpha_{n}=\sqrt{2}n^{-1/2}$ and $X_{n}\sim\mathcal{N}(0,1)$ for all $n\in\mathbb{N}$, their model coincides with the random Fourier series (1.1) with $a$ taken to be a Gaussian sequence, up to a scaling of space by $2\pi$.
@@
In [SW20], it is shown that a stochastic approximation of the logarithm of the Riemann zeta function may be decomposed into the sum of a Gaussian log-correlated field and a smooth field; using this, they show that the associated multiplicative chaos is absolutely continuous with respect to the GMC formed by the Gaussian part of the field with almost surely bounded Radon-Nikodym derivative. We emphasize that in our case, in contrast to [SW20], the Radon-Nikodym derivative of $\mu_{\gamma,a}$ with respect to the $\mu_{\gamma,g}$ constructed in Theorem 1.1 is not necessarily almost-surely bounded.
@@
Recently, [2024arXiv240817219V] has shown that one can reconstruct the underlying log-correlated field from its multiplicative chaos measure (at any inverse temperature less than or equal to critical) for what they call “mildly non-Gaussian” fields, that is, fields of the form $X=G+H$, with $G$ a log-correlated Gaussian field and $H$ a Hölder continuous field.
@@
The critical two-dimensional Stochastic Heat Flow (SHF) is a logarithmically-correlated process of random Borel measures on $\mathbb{R}^{2}$, constructed in [CSZ23flow], arising from directed polymers in random environment. In [CSZ23gmc], it is shown that the one-time marginal of the SHF is not a GMC. There, the question is raised “whether the critical 2d Stochastic Heat Flow is absolutely continuous w.r.t. some GMC.” We do not address this question in the present article.
@@
Most relevant to this article is [Junnila], which, as mentioned above, proves for a quite general notion of log-correlated fields $(X_{k}(t))_{k\in\mathbb{N},t\in[0,1]^{d}}$ on $[0,1]^{d}$ that, for any $K\subseteq U\subseteq\mathbb{C}$ with $K$ compact and $U\supset(0,\sqrt{2d})$ open, there is a $p>1$ for which the martingale
@@
converges in $L^{p}(\Omega)$ for all $\gamma\in K$ and $f\in C([0,1]^{d},\mathbb{C})$. To our knowledge, in this general case, independence of the limiting chaos measure $\mu_{\infty,\gamma}$ from the discrete approximation of $\sum_{k=1}^{\infty}X_{k}(t)$, the study of the multifractal spectrum, construction of the chaos at the critical $\gamma=\sqrt{2d}$, and more generally the question of which properties of GMC can be transfered to $\mu_{\infty,\gamma}$ are still open.

### v2

@@ Abstract.
In this article, we consider the multiplicative chaos measure associated to the log-correlated random Fourier series, or random wave model, with i.i.d. coefficients taken from a general class of distributions. This measure was shown to be non-degenerate when the inverse temperature is subcritical by Junnila (Int. Math. Res. Not. 2020 (2020), no. 19, 6169-6196). When the coefficients are Gaussian, this measure is an example of a Gaussian multiplicative chaos (GMC), a well-studied universal object in the study of log-correlated fields. In the case of non-Gaussian coefficients, the resulting chaos is not a GMC in general. However, we construct a coupling between the non-Gaussian multiplicative chaos measure and a GMC such that the two are almost surely mutually absolutely continuous.
@@ 1. Introduction
This article considers the multiplicative chaos measures associated to non-Gaussian log-correlated fields and their shared properties with respect to Gaussian multiplicative chaos (GMC), formalized through the relation of absolute continuity. The model we analyze is the log-correlated random Fourier series, or random wave model. We focus on the space $[0,1]$ for ease of exposition and address general bounded domains in $\mathbb{R}^{d}$, $d\geq 1$, in Section˜4.2. We begin by defining the random Fourier series and stating the main theorem. After this, we review GMC, discuss related literature in the non-Gaussian setting, and provide an outline of the proof of the main theorem.
@@
$\mu_{n,\gamma,a}(\mathrm{d}t):=\frac{e^{\gamma S_{n,a}(t)}}{Z_{n,\gamma,a}(t)}\,\mathrm{d}t\,$, (1.3)
where $Z_{n,\gamma,a}(t):=\mathbb{E}\left[e^{\gamma S_{n,a}(t)}\right]$ so that $\mu_{n,\gamma,a}[A]$ is a martingale with respect to the natural filtration of $a$ for all Borel sets $A$.
In [25], Junnila showed that for $\gamma\in(0,\sqrt{2})$ there exists a non-degenerate measure $\mu_{\gamma,a}$ such that $\mu_{n,\gamma,a}$ converges almost surely to $\mu_{\gamma,a}$ as $n\to\infty$ in the space of Radon measures on $[0,1]$ equipped with the topology of weak convergence (for $\gamma\geq\sqrt{2}$, the limiting measure is degenerate). We call $\mu_{\gamma,a}$ the multiplicative chaos measure associated to the fields $(S_{n,a})_{n\in\mathbb{N}}$.
When $a=g$ is a sequence of i.i.d. standard Gaussians, the non-degeneracy of $\mu_{\gamma,g}$ has been known since Kahane [27], and $\mu_{\gamma,g}$ is called Gaussian multiplicative chaos (GMC). GMC is a fractal random measure with fascinating properties and many important applications in modern probability; we review this object in Section 1.1 below. Here, we mention that GMC is a universal object expected to appear when exponentiating a field that exhibits asymptotically Gaussian behavior and covariance diverging logarithmically on short scales.
While the field $(S_{n,a}(t))_{t\in[0,1]}$ is log-correlated (shown in [25, Section 3.1]), the chaos measure $\mu_{\gamma,a}$ is not necessarily a GMC in general due to the contribution of, for instance, the first (non-Gaussian) term in the sum defining $S_{n,a}(t)$. This begs the question: for general sequences $a$, what properties does $\mu_{\gamma,a}$ share with GMC?
@@
Let $\gamma\in(1,\sqrt{2})$ and take $(a^{(i)}_{k})_{i\in\{1,2\},k\in\mathbb{N}}$ i.i.d. satisfying Eq.˜1.2. There exists a sequence of i.i.d. standard Gaussian random variables $g:=(g_{k}^{(i)})_{i\in\{1,2\},k\in\mathbb{N}}$ coupled to $a:=(a^{(i)}_{k})_{i\in\{1,2\},k\in\mathbb{N}}$ such that
$\mu_{\gamma,g}\ll\mu_{\gamma,a}\ll\mu_{\gamma,g}\qquad\text{almost surely.}$
The $(g^{(i)}_{k})_{k\in\mathbb{N},i\in\{1,2\}}$ are defined in (2.23), which makes use of (2.3), (2.19) and Lemma 2.6. We expect ˜1.1 can assist in proving a variety of almost-sure properties of both $\mu_{\gamma,a}$ and the associated approximating sequence of non-Gaussian fields $(S_{n,a})_{n\in\mathbb{N}}$ shared by the GMC $\mu_{\gamma,g}$ and the Gaussian log-correlated fields $(S_{n,g})_{n\in\mathbb{N}}$, for $\gamma\in(1,\sqrt{2})$.
Remark 1.1 (Higher dimensions).
In Theorem˜4.1, we extend our result to random Fourier series in bounded domains in $\mathbb{R}^{d}$.
Remark 1.2 (The result holds in the full subcritical regime on unit cubes).
Our proof uses as a blackbox a high-dimensional coupling result dating back to Yurinskii [45]. Sometime after the posting of this paper, a new high-dimensional coupling was developed in [14]. They verify that replacing the Yurinskii coupling with the coupling of [14] in our proof and then following the steps of our proof exactly yields ˜1.1 in the full subcritical regime $(0,\sqrt{2})$. In regards to higher dimensions (see Remark˜1.1), they show that for unit cube domains $D=(0,1]^{d}$, their coupling tool applies in the same way to yield Theorem˜4.1 in the full subcritical regime $(0,\sqrt{2d})$. In Section˜1.3, we discuss in more detail why our application of the Yurinskii coupling requires $\gamma>\sqrt{d}$. See Remark 4.2 for more on the restriction of [14] to unit cubes.
@@ 1.1. Background on GMC
Gaussian multiplicative chaos was introduced by Kahane in [27] as a mathematical model for energy dissipation in turbulence, making rigorous a program initiated by Mandelbrot in [33]. In [27], considering a locally compact metric space $(T,\rho)$ with reference (Radon) measure $\sigma$, Kahane gave a rigorous interpretation of the measure
@@
where $(X(x))_{x\in T}$ formally denotes a Gaussian field with $\sigma$-positive covariance kernel
@@
for $g$ bounded and continuous. The kernel $R$ is of $\sigma$-positive type if there exists a sequence $(R_{i})_{i}$ of continuous, pointwise nonnegative, and positive definite kernels $R_{i}:T^{2}\to\mathbb{R}_{+}$ with
@@
Formally, Eq.˜1.6 means that the Gaussian field $S$ can be approximated by the (pointwise-defined) Gaussian fields $S_{n}=\sum_{i=1}^{n}X_{i}$, where $X_{i}$ is an independent centered Gaussian with covariance $R_{i}$. The measure $M_{\gamma}$ in Eq.˜1.4 is then defined as the limit of the martingale sequence
$M_{n,\gamma}(\mathrm{d}x):=e^{\gamma S_{n}(x)-\frac{\gamma^{2}}{2}\operatorname{Var}[S_{n}(x)]}\ \sigma(\mathrm{d}x)\,.$
An important property proven in [27] is that the law of $M_{\gamma}$ is independent of the sequence $(S_{n})_{n\in\mathbb{N}}$ approximating $S$, so that we can associate a unique measure $M_{\gamma}$, called GMC, to the log-correlated Gaussian field $(X(x))_{x\in T}$. The later result of [42] shows that two approximating sequences yield the same $M_{\gamma}$ almost surely.
Removing the restriction Eq.˜1.6 was achieved in several steps. First, [40] defined $M_{\gamma}$ in the case that $R$ is a stationary positive definite kernel of the form Eq.˜1.5 using a mollification approach, which in particular does not rely on a martingale approximation. Later, [42] defined the GMC associated to general Gaussian fields using an approach based on random shifts. Finally, [6] simplified the mollification approach and gave a simple proof that the limiting measure is independent of the chosen mollification procedure. More on these approaches, as well as a rich theory on the behavior of GMC and its applications, can be found in [38, 7].
There are various examples of log-correlated fields that are non-Gaussian in the pre-limit producing a limiting GMC coming from random matrix theory; see, for instance, [44, 8, 13, 34, 15, 9, 28, 26, 31]. In [15], conditions on the exponential moments of the field under which one has a GMC limit are developed. These criteria and the fact that GMC is supported on its thick points have been used in [29, 30, 35] to determine the first order of the maximum of non-Gaussian log-correlated fields.
@@ 1.2. Results on non-Gaussian multiplicative chaos
$\mu(\mathrm{d}t)\stackrel{{\scriptstyle d}}{{=}}Z_{0}\mathbbm{1}_{[0,1/2]}(t)\mu_{0}(2\mathrm{d}t)+Z_{1}\mathbbm{1}_{[1/2,1]}(t)\mu_{1}(2\mathrm{d}t-1)\,$, (1.7)
@@
where $\mu_{0}$, $\mu_{1}$ are independent copies of $\mu$ and $Z_{0}$, $Z_{1}$ are nonnegative, non-Gaussian, i.i.d. with mean 1. Motivated by Eq.˜1.7, [3] studies multiplicative chaos measures for which the underlying field is given by the mass of certain cones under an infinitely divisible, independently scattered random measure. They show that, for their construction, a limiting measure exists and give a criterion for non-degeneracy and existence of moments. In [37], also motivated by Eq.˜1.7, so-called $\star$-scale invariant random measures are studied, which, under mild assumptions, are formally of the form $e^{\gamma L_{x}}\;\mathrm{d}x$ with $L$ a log-correlated and infinitely divisible process. They show the existence of $\star$-scale invariant random measures, characterize them (up to technical assumptions), and give a condition for non-degeneracy.
@@
Brownian multiplicative chaos is the chaos measure one obtains by formally exponentiating the square root of the occupation field of 2D Brownian motion. This object was constructed in [4] in the $L^{2}$-regime and in [1, 22] in the full subcritical regime. Great progress has been made in its study and characterization since then [2, 23, 24].
@@
In [18], random trigonometric series of the form $\sum_{n}\alpha_{n}X_{n}\cos(nt+\Phi_{n})$ are considered, where $\alpha_{n}\in\ell^{4}(\mathbb{R})\setminus\ell^{2}(\mathbb{R})$, $\Phi_{n}$ are i.i.d. with $\Phi_{1}\sim\mathrm{Unif}([0,2\pi))$, and $X_{n}$ are independent with $\mathbb{E}[X_{n}^{2}]=1$ and $\mathbb{E}[X_{n}^{2m}]=O(K^{m}m!)$ for some $K>0$ and all $m\geq 2$, $n\geq 1$. They study the associated chaos measures and give results on the Hausdorff dimension. It is worth mentioning that in the case $\alpha_{n}=\sqrt{2}n^{-1/2}$ and $X_{n}\sim\mathcal{N}(0,1)$ for all $n\in\mathbb{N}$, their model coincides with the random Fourier series (1.1) with $a$ taken to be a Gaussian sequence, up to a scaling of space by $2\pi$.
@@
In [41], it is shown that a stochastic approximation of the logarithm of the Riemann zeta function may be decomposed into the sum of a Gaussian log-correlated field and a smooth field; using this, they show that the associated multiplicative chaos is absolutely continuous with respect to the GMC formed by the Gaussian part of the field with almost surely bounded Radon-Nikodym derivative. We emphasize that in our case, in contrast to [41], the Radon-Nikodym derivative of $\mu_{\gamma,a}$ with respect to the $\mu_{\gamma,g}$ constructed in Theorem 1.1 is not necessarily almost surely bounded.
A recent series of papers [19, 20, 21] constructs a non-Gaussian multiplicative chaos measure, in the subcritical and critical regimes, associated to a Dirichlet series of coefficients coming from a certain class of random multiplicative functions restricted to the critical line. In fact, their results are robust under small perturbations away from the critical line. They do not attempt to obtain a coupling between their chaos measures and GMC measures. Relatedly, their methods do not involve Gaussian approximation.
@@
Recently, [43] has shown that one can reconstruct the underlying log-correlated field from its multiplicative chaos measure (at any inverse temperature less than or equal to critical) for what they call “mildly non-Gaussian” fields, that is, fields of the form $X=G+H$, with $G$ a log-correlated Gaussian field and $H$ a Hölder continuous field.
@@
The critical two-dimensional Stochastic Heat Flow (SHF) is a logarithmically-correlated process of random Borel measures on $\mathbb{R}^{2}$, constructed in [10], arising from directed polymers in random environment. In [11], it is shown that the one-time marginal of the SHF is not a GMC. There, the question is raised “whether the critical 2d Stochastic Heat Flow is absolutely continuous w.r.t. some GMC.” We do not address this question in the present article, however, after our paper was posted, a “conditional GMC” structure was uncovered in [16].
@@
Most relevant to this article is [25], which, as mentioned above, proves for a quite general notion of log-correlated fields $(X_{k}(t))_{k\in\mathbb{N},t\in[0,1]^{d}}$ on $0,1]^{d}$ that, for any $K\subseteq U\subseteq\mathbb{C}$ with $K$ compact and $U\supset(0,\sqrt{2d})$ open, there is a $p>1$ for which the martingale
@@
converges in $L^{p}(\Omega)$ for all $\gamma\in K$ and $f\in C([0,1]^{d},\mathbb{C})$. To our knowledge, in this general case, independence of the limiting chaos measure $\mu_{\gamma}$ from the discrete approximation of $\sum_{k=1}^{\infty}X_{k}(t)$, the study of the multifractal spectrum, construction of the chaos at the critical $\gamma=\sqrt{2d}$, and more generally the question of which properties of GMC can be transfered to $\mu_{\gamma}$ are still open.